OpenStudy (anonymous):

Solve the equation. –2 3/7 + b = 6 1/7 A. b = 3 5/7 B. b = 4 2/7 C. b = 8 2/7 D. b = 8 4/7

OpenStudy (anonymous):

I think it might be B.

OpenStudy (k12awesomeness):

I know the answer

OpenStudy (anonymous):

Well first take - 2 3/7 - 6 1/7

OpenStudy (k12awesomeness):

The answers D.

OpenStudy (astrophysics):

\[-2 \frac{ 3 }{ 7 } +b = 6 \frac{ 1 }{ 7 } \implies b = 6\frac{ 1 }{ 7 } +2\frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

The answer is C. 3/7 - 1/7 = 2/7

OpenStudy (anonymous):

8 2/7

OpenStudy (k12awesomeness):

@kittenlover01 the answer is D.

OpenStudy (astrophysics):

You should try to teach the user rather than giving a direct answer.

OpenStudy (k12awesomeness):

yeah yeah... my explaining is hard to understand, so I didn't want to confuse her, I gaveher the answer.

OpenStudy (anonymous):

ok thank you all

OpenStudy (k12awesomeness):

No Problem!! :)

OpenStudy (mathmale):

To all of you who have given direct answers such as "the answer is D," kindly stop; you are violating OpenStudy's code of conduct by doing so. mathmale, Moderator